# Necessary and sufficient conditions for correctness of complex Langevin

### v1

@@
( August 20, 2025)
We derive a family of correctness conditions for complex Langevin simulations. In particular, we show that if in a given theory the expectation values of all observables within a particular space satisfy the theory’s Schwinger–Dyson equations as well as certain bounds, then these expectation values are necessarily correct. In fact, these findings are not only valid in the context of complex Langevin simulations, but they also hold for general probability densities on complex manifolds, given an initial complex density on a real manifold. We test these criteria in a few simple one- and two-dimensional toy models and find that they are indeed capable of ruling out incorrect results without the need of exact solutions.
@@ 1 Introduction
Establishing the correctness of complex Langevin simulation results is thus of utmost importance in order to ensure the reliability of the method. While there is a plethora of potential correctness criteria to be found in the literature, none of them appear to be entirely satisfactory and free of loopholes. In this work, we attempt to fill this gap by deriving a set of conditions that are both necessary and sufficient to ensure that a probability density $P$ on $\mathcal{M}_{c}$ produces the same expectation values as a given complex density $\rho$ on $\mathcal{M}_{r}$. The correctness criterion proposed here can thus be applied to complex Langevin results, but it is, in fact, more general than that. Indeed, it can be used in any situation where a complex density on a real manifold $\mathcal{M}_{r}$ is traded for a probability density on the complexification of $\mathcal{M}_{r}$.
@@ 2.1 Manifolds and densities
This is not really a restriction because by Cauchy’s theorem we can always deform or shift the integration manifold $\mathcal{M}_{r}$ away from any zeroes without changing the expectation values of observables.
@@ 2.2 Observables and Schwinger–Dyson operators
It is not hard to see that all $f\in\mathcal{H}$ are integrable with respect to $|\rho|d\mu$.
@@ 2.3 Correctness of probability measures on $\mathcal{M}_{c}$
To be clear, since we assume the partial derivatives of $\rho$ to grow at most polynomially, these conditions require that $w(\mathbf{x})$ decays faster than any power of $|\mathbf{x}|$, while (20) must hold at the same time. In fact, for reasons stated below, we actually require $w(\mathbf{x})$ to decay at least like a Gaussian, i.e.,
@@ 3.2 Sufficiency of the necessary conditions
Note that $\tilde{\mathcal{H}}\subset\mathfrak{C}_{0}(\mathcal{M}_{r})$, where $\mathfrak{C}_{0}(\mathcal{M}_{r})$ is the space of all continuous functions on $\mathcal{M}_{r}$ vanishing at infinity. This is due to our requirements that $\rho$ does not vanish on $\mathcal{M}_{r}$ and that any possible large growth of $f$ at infinity is counteracted by the fast decay of $w$. For the latter to be true, it was necessary to restrict the growth of the partial derivatives of $\rho$ earlier.
@@
Any bounded linear functional $\tilde{T}$ on $\mathfrak{C}_{0}(\mathcal{M}_{r})$ is given by a complex measure, which we represent by a density $\tilde{\sigma}$ on $\mathcal{M}_{r}$ as follows:
@@
Restricting to the subset $\tilde{\mathcal{H}}\subset\mathfrak{C}_{0}(\mathcal{M}_{r})$ and applying this proposition to the functional $\tilde{T}_{P}$ generated by the probability measure $dP$ according to (30) and (29), we find that it is represented by a complex measure $\tilde{\sigma}_{P}$ of bounded total variation, i.e.,
@@
Note that under our conditions, there are once again no boundary terms here. Up to now, we have treated $\sigma_{P}$ as a distribution, representing a measure, but we have not yet addressed the question whether it is differentiable, so derivatives have to be understood in the distributional sense.
@@
Indeed, this will turn out to be true because there are enough functions in $\mathfrak{A}$ such that a distribution $T$ satisfying $(T,f)=0\;\forall f\in\mathfrak{A}$ vanishes identically. However, in order to see this more explicitly, we require a few more arguments.
@@
where $h$ is in the subalgebra $\mathfrak{A}_{G}$ of $\mathfrak{A}$ consisting of functions that do not depend on $\mathbf{x}$. Note that the functions $f_{\mathbf{k}}$ are not in $\mathfrak{A}$, but they are contained in its closure $\overline{\mathfrak{A}}$ with respect to $\|\cdot\|_{w}$, as is shown in Appendix B.
@@
respectively. It should be understood that the practical applicability of the proposed correctness criterion might strongly depend on all of these choices.
@@ 5 Examples and numerical checks
In this section, we perform numerical tests on the new condition for correctness for a handful of simple systems with known results. We restrict ourselves to models with $m=0$, i.e., there are no Lie-group-valued degrees of freedom, such that $\mathcal{M}_{r}=\mathbb{R}^{n}$. Following the FAIR guiding principles [43], we make the simulation data underlying the results presented in this section [44], as well as the used analysis scripts [45], available online. Moreover, our simulation code is available upon request.
@@ 5.1.1 One dimension
The main message of Fig. 1 (left) is that there are four different regions in $m_{H}$ for which the SDEs are well satisfied and boundary terms vanish. Nonetheless, it is known from [46] that only one of these regions corresponds to correct solutions. Thus, we are indeed confronted with a situation where the study of boundary terms (or SDEs) alone is not sufficient to make definite statements about correct convergence. Let us therefore invoke the new correctness criterion.
@@ 5.1.2 Two dimensions
We have thus established that the new correctness criterion can detect incorrect convergence due to unwanted integration cycles in simple quartic models in one and two dimensions. We believe that these results can be extended to quartic models in higher dimensions without too many complications. Regarding models with more involved polynomial actions, we are confident that one should be able to find viable control observables $f(\mathbf{z})$ reasonably easily, but we have not attempted to do so yet. At the time of writing, we are not aware of any good guiding principle on how to choose appropriate $f(\mathbf{z})$ for general theories. This will become particularly clear for the model studied in the next subsection.
@@ 5.2 Densities with zeros
It was obtained by minimizing the distance [^5] One possible approach to find suitable polynomials in practice could thus be to start with an arbitrary function for which one can somehow show that it violates the bounds (26) and then approximate it (in the norm (19)) with a polynomial of sufficiently high order. We also remark that we are, in fact, not restricted to using polynomials here. Any element of $\mathcal{H}$ would work in principle. to the function $e^{-2(1+(2\mathrm{i}+z))^{2}}$, which we found (rather by coincidence) to violate the bounds, within the norm (19). For this polynomial, we find that the complex Langevin evolution produces an expectation value whose absolute value reads $\left|\langle f\rangle_{CL}\right|=902$ with negligible statistical uncertainties, while the bound on the right-hand side of (26) reads $C\|f\|_{w}\approx 702$. This bound is thus clearly violated, from which one can safely conclude that the complex Langevin results are incorrect.
@@ 6 Towards more realistic models
$\frac{1}{\left\langle\frac{\rho^{\prime}(\mathbf{x},\mathbf{g})}{|\rho_{r}^{\prime}(\mathbf{x},\mathbf{g})|}\right\rangle_{|\rho_{r}^{\prime}|}}=\frac{1}{\left\langle e^{i\phi_{r}^{\prime}(\mathbf{x},\mathbf{g})}w(\mathbf{x})\right\rangle_{|\rho_{r}^{\prime}|}}\;.$ (94)
@@ 7 Summary & conclusions
More precisely, the necessary and sufficient conditions are on the one hand the validity of the Schwinger–Dyson equations (24) and (25), and on the other hand the validity of the bounds (26) for all $f\in\mathcal{H}$, which need to hold simultaneously. These results were derived using standard theorems of functional analysis, assuming only that $\rho$ is holomorphic and normalizable and that the partial derivatives of $\rho$ (with respect to both $\mathbf{x}$ and $\mathbf{g}$) grow at most like a polynomial on $\mathcal{M}_{r}$. None of these assumptions is particularly restrictive in the view of usual theories of interest in physics. Moreover, we have required $\mathcal{M}_{r}$ to be free of any zeros of $\rho$, which, however, can always be achieved by a continuous deformation of $\mathcal{M}_{r}$. In particular, the presence of a fermionic determinant in the density, as it appears in the path integral treatment of quantum field theories with fermions, does not spoil the proof. Finally, the quantities $w$ and $\rho_{r}$, defined in (18), are to be chosen such that the former obeys (21) and that the constant $C$ in (20) as well as the norms $\|f\|_{w}$ in (19) (for all $f\in\mathcal{H}$) are finite.
@@ Acknowledgments
This research was funded in part by the Austrian Science Fund (FWF) via the Principal Investigator Project P36875. The numerical simulations, from which the results discussed in this work have been computed, were performed on the computing cluster of the University of Graz (GSC). The data analyses are based on the Python scientific computing ecosystem [49], in particular via the packages [50, 51, 52] and we are also grateful for the creation and maintenance of the dependencies of the latter. Moreover, auxiliary numerical calculations were performed with the Wolfram Mathematica computer algebra system [53].
@@
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any author-accepted manuscript version arising.

### v2

@@
( December 10, 2025)
We derive a family of correctness conditions for complex Langevin simulations. In particular, we show that if in a given theory the expectation values of all observables within a particular space satisfy the theory’s Schwinger–Dyson equations as well as certain bounds, then these expectation values are necessarily correct. In fact, these findings are not only valid in the context of complex Langevin simulations, but they also hold for general probability densities on complex manifolds, given an initial complex density on a real manifold. We stress that, while the proposed conditions are necessary and sufficient in a mathematical sense, their practical use is not to prove the correctness of obtained simulation results. Rather, they are mainly useful for detecting incorrect convergence. In particular, we test these criteria in a few simple one- and two-dimensional toy models and find that they are indeed capable of ruling out incorrect results without the need of exact solutions.
@@ 1 Introduction
Establishing the correctness of complex Langevin simulation results is thus of utmost importance in order to ensure the reliability of the method. While there is a plethora of potential correctness criteria to be found in the literature, as we shall discuss, none of them appear to be entirely satisfactory and free of loopholes. In this work, we attempt to fill this gap by deriving a set of conditions that are both necessary and sufficient to ensure that a probability density $P$ on $\mathcal{M}_{c}$ produces the same expectation values as a given complex density $\rho$ on $\mathcal{M}_{r}$. The correctness criterion proposed here can thus be applied to complex Langevin results, but it is, in fact, more general than that. Indeed, it can be used in any situation where a complex density on a real manifold $\mathcal{M}_{r}$ is traded for a probability density on the complexification of $\mathcal{M}_{r}$, see, e.g., [4, 5, 6, 7].
@@ 2.1 Manifolds and densities
This is not really a restriction because by Cauchy’s theorem we can always deform the integration manifold $\mathcal{M}_{r}$ away from any zeros, for instance by shifting by some vector with small purely imaginary components, without changing the expectation values of observables.
@@ 2.2 Observables and Schwinger–Dyson operators
It is not hard to see that all $f\in\mathcal{H}$ are integrable with respect to $|\rho|d\mu$. In Section 5.2, we shall discuss a specific example in which $\mathfrak{A}\neq\mathcal{H}$.
@@ 2.3 Correctness of probability measures on $\mathcal{M}_{c}$
To be clear, since we assume the partial derivatives of $\log\rho$ to grow at most polynomially, these conditions require that $w(\mathbf{x})$ decays faster than any power of $|\mathbf{x}|$, while (20) must hold at the same time. In fact, for reasons stated below, we actually require $w(\mathbf{x})$ to decay at least like a Gaussian, i.e.,
@@ 3.2 Sufficiency of the necessary conditions
Note that $\tilde{\mathcal{H}}\subset\mathfrak{C}_{0}(\mathcal{M}_{r})$, where $\mathfrak{C}_{0}(\mathcal{M}_{r})$ is the space of all continuous functions on $\mathcal{M}_{r}$ vanishing at infinity. This is due to our requirements that $\rho$ does not vanish on $\mathcal{M}_{r}$ and that any possible large growth of $f$ at infinity is counteracted by the fast decay of $w$. For the latter to be true, it was necessary to restrict the growth of the partial derivatives of $\log\rho$ earlier.
@@
Any bounded linear functional $\tilde{T}$ on $\mathfrak{C}_{0}(\mathcal{M}_{r})$ is given by a complex measure, which we represent by a distributional density $\tilde{\sigma}$ on $\mathcal{M}_{r}$ as follows:
@@
Restricting to the subset $\tilde{\mathcal{H}}\subset\mathfrak{C}_{0}(\mathcal{M}_{r})$ and applying this proposition to the functional $\tilde{T}_{P}$ that is generated by the probability measure $dP$ according to (30) and (29) and that we have just shown to be bounded, we find that it is represented by a complex measure $\tilde{\sigma}_{P}$ of bounded total variation, i.e.,
@@
Note that under our conditions, there are once again no boundary terms here. Up to now, we have treated $\sigma_{P}$ as a distribution, in fact one of order zero [10] according to the bounds (26). $\sigma_{P}$ therefore represents a complex (Borel) measure by the Riesz-Markov theorem. But we have not yet addressed the question whether it is a differentiable function, so the derivatives in (40) and (41) have to be understood in the distributional sense.
@@
These differential equations then would imply that $\sigma_{P}$ is really a differentiable function. Indeed, this will turn out to be true because there are enough functions in $\mathfrak{A}$ such that a distribution $T$ satisfying $(T,f)=0\;\forall f\in\mathfrak{A}$ vanishes identically. However, in order to see this more explicitly, we require a few more arguments.
@@
where $h$ is in the sub-algebra $\mathfrak{A}_{G}$ of $\mathfrak{A}$ consisting of functions that do not depend on $\mathbf{x}$. Note that the functions $f_{\mathbf{k}}$ are not in $\mathfrak{A}$, but they are contained in its closure $\overline{\mathfrak{A}}$ with respect to $\|\cdot\|_{w}$, as is shown in Appendix B.
@@
respectively. It should be understood that the practical applicability of the proposed correctness criterion might strongly depend on all of these choices. A detailed analysis systematically comparing different possible correctness conditions arising from our proposal is, however, beyond the scope of this work.
@@ 5 Examples and numerical checks
In this section, we perform numerical tests on the new condition for correctness for a handful of simple systems with known results. We restrict ourselves to models with $m=0$, i.e., there are no Lie-group-valued degrees of freedom, such that $\mathcal{M}_{r}=\mathbb{R}^{n}$. A more detailed analysis in the case of gauge theories is deferred to future work. Following the FAIR guiding principles [45], we make the simulation data underlying the results presented in this section [46], as well as the used analysis scripts [47], available online. Moreover, our simulation code is available upon request.
@@ 5.1.1 One dimension
The main message of Fig. 1 (left) is that there are four different regions in $m_{H}$ for which the SDEs are well satisfied and boundary terms vanish. Nonetheless, it is known from [48] that only one of these regions corresponds to correct solutions. Thus, we are indeed confronted with a situation where the study of boundary terms (or SDEs) alone is not sufficient to make definite statements about correct convergence. This was found to be the case also in another simple model in [50]. The same is true for the widely-employed correctness criterion based on the decay properties of the complex-Langevin drift term [33]. Indeed, as was shown in [43], the respective distributions of $z$ in the complex plane for, e.g., $m_{H}=10$ and $m_{H}=34$ are identical up to a rotation. This can be understood from the fact that for both these choices of kernel the complex Langevin evolution is equivalent to a real Langevin evolution, albeit on a rotated real axis, the rotation angle being the only thing distinguishing between them. As a consequence, the distributions of drift magnitudes must be equivalent by definition. Clearly, thus, the drift criterion cannot detect this incorrect convergence. Let us therefore invoke the new correctness criterion.
@@ 5.1.2 Two dimensions
We have thus established that the new correctness criterion can detect incorrect convergence due to unwanted integration cycles in simple quartic models in one and two dimensions. We believe that these results can be extended to quartic models in higher dimensions without too many complications. Regarding models with more involved polynomial actions, we are confident that one should be able to find viable control observables $f(\mathbf{z})$ reasonably easily, but we have not attempted to do so yet. We stress that $f(z)$ and $f(z_{1},z_{2})$ above were chosen in a purely ad-hoc manner and not derived systematically. At the time of writing, we are not aware of any good guiding principle on how to choose appropriate $f(\mathbf{z})$ for general theories. This will become particularly clear for the model studied in the next subsection.
@@ 5.2 Densities with zeros
It was obtained by minimizing the distance [^5] One possible approach to find suitable polynomials in practice could thus be to start with an arbitrary function for which one can somehow show that it violates the bounds (26) and then approximate it (in the norm (19)) with a polynomial of sufficiently high order. We also remark that we are, in fact, not restricted to using polynomials here. Any element of $\mathcal{H}$ would work in principle. $\|f(z)-g(z)\|_{w}$ to the function $g(z)=e^{-2(1+(2\mathrm{i}+z)^{2})}$, which we found (rather by coincidence) to violate the bounds, within the norm (19). For this polynomial, we find that the complex Langevin evolution produces an expectation value whose absolute value reads $\left|\langle f\rangle_{CL}\right|=902$ with negligible statistical uncertainties, while the bound on the right-hand side of (26) reads $C\|f\|_{w}\approx 702$. This bound is thus clearly violated, from which one can safely conclude that the complex Langevin results are incorrect.
@@ 6 Towards more realistic models
$\frac{1}{\left\langle\frac{\rho^{\prime}(\mathbf{x},\mathbf{g})}{|\rho_{r}^{\prime}(\mathbf{x},\mathbf{g})|}\right\rangle_{|\rho_{r}^{\prime}|}}=\frac{1}{\left\langle e^{i\varphi_{r}^{\prime}(\mathbf{x},\mathbf{g})}w(\mathbf{x})\right\rangle_{|\rho_{r}^{\prime}|}}\;.$ (94)
@@ 7 Summary & conclusions
More precisely, the necessary and sufficient conditions are on the one hand the validity of the Schwinger–Dyson equations (24) and (25), and on the other hand the validity of the bounds (26) for all $f\in\mathcal{H}$, which need to hold simultaneously. These results were derived using standard theorems of functional analysis, assuming only that $\rho$ is holomorphic and normalizable and that the partial derivatives of $\log\rho$ (with respect to both $\mathbf{x}$ and $\mathbf{g}$) grow at most like a polynomial on $\mathcal{M}_{r}$. None of these assumptions is particularly restrictive in the view of usual theories of interest in physics. Moreover, we have required $\mathcal{M}_{r}$ to be free of any zeros of $\rho$, which, however, can always be achieved by a continuous deformation of $\mathcal{M}_{r}$. In particular, the presence of a fermionic determinant in the density, as it appears in the path integral treatment of quantum field theories with fermions, does not spoil the proof. Finally, the quantities $w$ and $\rho_{r}$, defined in (18), are to be chosen such that the former obeys (21) and that the constant $C$ in (20) as well as the norms $\|f\|_{w}$ in (19) (for all $f\in\mathcal{H}$) are finite.
@@ Acknowledgments
The authors thank Lorenzo Salcedo for useful comments on the manuscript. This research was funded in whole, or in part, by the Austrian Science Fund (FWF) [10.55776/P36875]. The numerical simulations, from which the results discussed in this work have been computed, were performed on the computing cluster of the University of Graz (GSC). The data analyses are based on the Python scientific computing ecosystem [52], in particular via the packages [53, 54, 55] and we are also grateful for the creation and maintenance of the dependencies of the latter. Moreover, auxiliary numerical calculations were performed with the Wolfram Mathematica computer algebra system [56].
@@
For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.